Two pumps are used to fill a water storage tank at a resort. One pump can fill the tank by itself in 9 hours, and the other can fill it in 6 hours. How long will it take both pumps operating together to fill the tank?
Answer by stanbon(75887)

Two pumps are used to fill a water storage tank at a resort. One pump can fill the tank by itself in 9 hours, and the other can fill it in 6 hours. How long will it take both pumps operating together to fill the tank?
-------------------------------------------------------------------
1st pump DATA:
time = 9 hr/job ; rate = (1/9) job/hr
-----------------------------------------
2nd pump DATA:
time = 6 hr/job ; rate = 1/6 job/hr
------------------------------------------
Together DATA:
time = x hr/job ; rate = 1/x job/hr
-------------------------------------------
Equation:
rate + rate = together rate
(1/9) + (1/6) = 1/x
2x + 3x = 18
5x = 18
x = 18/5 = 3.6 hrs (time required when both pumps are operating)
======================================================================
Cheers,
Stan H.